# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ V-D Factors Influencing LLMs Performance
Despite the remarkable capabilities of LLMs in code analysis tasks, their effectiveness is influenced by various factors, including the prompts used (further explored in Section VIII), and the state of the code being analyzed.
@@
One of the issues arises when resolved vulnerabilities remain in the code, often in the form of comments and annotations left by the developers. LLMs may mistakenly interpret these resolved issues as existing vulnerabilities, which can lower their ability to detect existing security vulnerabilities. This implies that there is a need for developers to standardize code commenting practices, to avoid confusion and improve LLMs performance in detecting security vulnerabilities. Additionally, when handling longer pieces of code, LLMs often struggle to identify smaller, more subtle security defects [54].
@@
Several other challenges affect LLMs performance in detecting vulnerabilities. These include difficulty in understanding developers’ intent, misinterpretation of code comments, and inability to execute code. Chen et al. [61] found that these challenges contributed to the high rate of false positives obtained when experimenting with ChatGPT as a vulnerability detection tool for smart contracts.
@@
Some authors have focused their research on identifying and addressing the factors that influence LLMs performance in code analysis tasks, aiming to enhance their accuracy and reliability. Specifically, Wang et al. [89] investigated the impact of naming on LLMs code analysis. In their experiment, they created datasets containing code with nonsensical or misleading variables, methods, and function names. Using CodeBERT for the analysis, they demonstrated that naming significantly affects LLMs ability to detect vulnerabilities, indicating that LLMs rely on clear and meaningful names.
@@
Fang et al. [90] identified a gap in research, concerning how effective LLMs are at analyzing confusing, obfuscated, or poorly readable code. The authors found that GPT models were more likely than smaller models to generate accurate and detailed explanations for input code. On the other hand, obfuscated code impacted the ability of LLMs to generate explanations. GPT models had a drop in security analysis accuracy, while smaller models were unable to handle obfuscated code. In conclusion, GPT models excel in code analysis but struggle with obfuscated code, indicating that clarity of the code significantly impacts LLMs analysis performance.
V-E LLMs Fine-tuning for Vulnerability Detection
@@
Figure 3: Fine-tuning LLM with the security-specific dataset.
@@
In their work, Shestov et al. [91] fine-tuned the LLM model WizardCoder [92] to improve it in vulnerability detection tasks for Java code. Their first contribution involved modifying the training process, which resulted in faster training times without compromising the model performance. The fine-tuned WizardCoder outperformed models such as CodeBERT, demonstrating the potential of fine-tuning LLMs for specialized source code analysis tasks. The authors also noted that vulnerability detection faces a fundamental challenge of class imbalance, with far fewer positive cases (i.e., the presence of vulnerabilities) than negative cases. By applying a combination of focal loss and sample weighting, the authors showed that models can achieve notable performance, even when dealing with such imbalances.
@@
After evaluating LLMs for vulnerability detection, the authors of [57] fine-tuned one of the models that performed poorly on vulnerability detection tasks, Meta-Llama-3-8B-Intruct, to verify the impact of fine-tuning on the domain-specific performance. The results showed that fine-tuning can have a positive impact on vulnerability detection tasks, but challenges in identifying exact vulnerabilities remain. Moreover, fine-tuning was also explored by Ding et al. [58]. Despite applying advanced techniques, such as class weights and contrastive learning, LLMs struggled with detecting realistic vulnerabilities.
@@
Zhang et al. [46] introduced two fine-tuning approaches for vulnerability localization: discriminative fine-tuning and generative fine-tuning. The discriminative fine-tuning approach treats vulnerability localization as a sequence labeling task, classifying each code statement as either vulnerable or not. In contrast, generative fine-tuning trains models to generate both the line numbers and the corresponding vulnerable statements within a function. Currently, although generative fine-tuning appears promising, it is outperformed by discriminative fine-tuning, which significantly enhances model performance.
@@
Alqarni et al. [93] conducted experimental research on BERT with the aim to improve its ability on vulnerability detection tasks. Although BERT was not originally designed for this domain, they improved its performance by integrating additional layers into a new model, and retraining it using a GitHub vulnerability dataset. The BERT base models (two versions) use token, segment, and position embeddings to understand context, masking 15% of tokens to predict them during training. This new model, which integrates with BERT and shares both embeddings and attention mechanisms, achieved a notable test accuracy of 99.30%.
@@
Basharat et al. [94] developed a method for detecting vulnerabilities in code using LLMs, by combining pattern-exploiting training and cloze-style queries to fine-tune models. Following the model of cloze tests, they hid parts of the code and asked the model to predict the missing parts. By performing this task on datasets containing vulnerable code, the model learns to recognize patterns that indicate potential security weaknesses. Results indicate that the method, especially using the RoBERTa model, can effectively identify vulnerabilities in different programming languages, such as C/C++.
@@
Guo et al. [95] compared 6 LLMs trained for vulnerability detection against 6 general-purpose LLMs, 3 of which were further fine-tuned on a dataset they compiled. Their findings demonstrate that fine-tuning significantly improves the performance of smaller models, enabling them to outperform larger, general-purpose models in specific contexts. Fine-tuned models achieved remarkable results, with CodeLlama-7b-fine-tuned obtaining an F1-score of 97%, compared to the best performance of around 51% obtained by non-fine-tuned models. The study revealed that, while fine-tuning smaller models can improve their performance, they are limited to the exact tasks for which they were trained. Additionally, the research underscores issues with current datasets, including labeling errors, which can adversely affect model training and performance.
VI LLMs for Vulnerability Fixes
@@
Developers often introduce security vulnerabilities in their code, which, if exploited, can compromise system and application security. Addressing these vulnerabilities presents significant challenges, as manual mitigation is time-consuming, prone to errors, and requires experienced developers with specialized knowledge. LLMs can assist in identifying the root causes of these issues and providing the right fixes, significantly reducing developers’ workloads [96, 97]. This section explores the usage of LLMs for patching vulnerabilities, as well as the possibility of improving this usage through fine-tuning processes.
VI-A Uses of LLMs for Fixing Vulnerabilities
@@
Several studies focused on LLMs abilities to fix security vulnerabilities in code. For instance, Nong et al. [98] used the power of LLMs for vulnerability patching without needing specific test inputs or training/fine-tuning. To achieve this, they developed LLMPatch, an automated LLM-based patching system. They evaluated LLMPatch on real-world vulnerable code, with a focus on zero-day vulnerabilities. LLMPatch was able to patch vulnerabilities and successfully surpassed traditional patching strategies based on DL (VulRepair [99]) and abstract syntax trees (Getafix [100]). Moreover, LLMPatch proved to be cost-effective and time-efficient.
@@
Berabi et al. [101] proposed a new method for fixing code vulnerabilities using LLMs, which enhances their ability to fix vulnerabilities. To increase the model performance, their approach combines static code analysis with code reduction, a technique that involves selecting only the code that contains a defect and the relevant context for interpreting it, and feeding both to the LLM. Results showed that this approach improved the efficiency of the model, by allowing it to understand and focus only on the sections that require fixing.
@@
Zhang et al. [97] demonstrated the ability of ChatGPT-4 and Claude to fix memory corruption vulnerabilities in real-world C/C++ code. LLMs can effectively address localized issues, such as memory leaks (CWE-401), but struggle with complex issues involving external dependencies, precise context, and intricate code structures. Furthermore, the authors highlighted that providing more context, such as external functions, data structures, and explanations of code logic, showed a positive effect on LLMs performance.
@@
Pearce et al. [96] explored how well LLMs, such as OpenAI’s Codex and AI21’s Jurassic J-1, can fix software vulnerabilities in a zero-shot context. They evaluated the performance of 5 commercially available, black-box, “off-the-shelf” LLMs, as well as an open-source model, and their own locally-trained model on their dataset. They discovered that, while these models could successfully repair synthetic and hand-crafted vulnerabilities, they faced challenges with real-world scenarios, due to the complexity of prompt design and natural language variations.
VI-B LLMs Fine-tuning for Fixing Vulnerabilities
Several researchers explored fine-tuning as a way to enhance LLMs effectiveness in fixing vulnerabilities.
@@ VI-B LLMs Fine-tuning for Fixing Vulnerabilities
De et al. [102] fine-tuned two LLMs, Code Llama and Mistral, and compared them with traditional DL-based approaches, including VulRepair, VRepair [103], and VulMaster [104]. The study also highlighted the problem of data overlap in previous research, revealing that models trained on overlapping datasets showed inflated performance metrics. To address this, their evaluation involved two datasets: an original dataset, used in other papers, and a refined dataset, which was curated by the authors to avoid overlaps between training and test data. This approach helped ensure a more accurate assessment of the model performance on new, unseen data. The results showed that the proposed method offered significant improvements in terms of code modifications efficiency and accuracy, compared to the DL-based methods.
@@
Wu et al. [105] evaluated the performance of different LLMs in fixing Java vulnerabilities, with and without fine-tuning, on two benchmarks they proposed. Across 5 non-fine-tuned LLMs, the best-performing model, Codex, fixed 20.4% of vulnerabilities. Concerning the fine-tuning experiment, the authors could fine-tune 4 models out of 5, since OpenAI does not provide a public API for fine-tuning Codex. Among the fine-tuned models, the best performer (i.e., fine-tuned InCoder), fixed only 18% of vulnerabilities, a small improvement of 8% over the non-fine-tuned version. Besides, these 4 fine-tuned models could only fix vulnerabilities that ureqired simple changes, such as deleting statements, renaming variables, or renaming methods. The models struggled with more complex vulnerabilities, such as CWE-325: Missing Cryptographic Step, and CWE-444: HTTP Request/Response Smuggling.
VII LLMs for Detecting and Fixing Vulnerabilities in a Unified Process
@@
Ensuring the security of increasingly complex software systems requires novel approaches that can keep pace with evolving threats. LLMs represent a promising tool for detecting and fixing software vulnerabilities [8].
@@
In Section V and Section VI, we discussed using LLMs as tools for handling vulnerability detection and vulnerability fixing separately. In this section, we explore an integrated approach where LLMs are utilized to detect and fix security vulnerabilities in a unified process. Additionally, it examines the capability of LLMs to identify vulnerabilities in the code they generate, and automatically resolve them. Last, the section covers fine-tuning for enhancing LLMs performance, in a unified pipeline for vulnerability detection and fixing.
VII-A LLMs for Detecting and Fixing Vulnerabilities
@@
Several studies have focused on both detecting and fixing software vulnerabilities using LLMs. For instance, Noever et al. [69] evaluated the capability of GPT-4 to detect software vulnerabilities and suggest fixes for such vulnerabilities. Their evaluation included 29 code samples from numerous repositories, including those from NASA and the US Department of Defense. GPT-4 demonstrated a strong ability to detect and fix vulnerabilities, achieving a 90% reduction in vulnerabilities with only an 11% increase in lines of code, while maintaining low rates of false positives.
@@
Le et al. [8] investigated the efficiency of ChatGPT and Bard in the detection and fixing of security vulnerabilities, focusing on JavaScript code. The results showed that ChatGPT outperformed Bard with a marginal higher accuracy of 71.66%, versus 68.33%.
VII-B LLMs for Detecting and Fixing Vulnerabilities in Self-produced Code
@@
Code generated by LLMs can contain vulnerabilities, as we have previously discussed in Section IV. However, researchers have shown that it is possible to re-prompt LLMs to identify and fix those vulnerabilities. By doing that, they have explored the self-evaluation approach and concluded that the code can be significantly improved through iterative refinement.
@@
The research conducted by Liu et al. [23] has shown promising results in addressing security vulnerabilities using ChatGPT, with over 89% of vulnerabilities successfully mitigated during a multi-round fixing process. In this approach, Liu et al. interacted with ChatGPT over several rounds, giving feedback on the provided code, and asking for improvements until the vulnerabilities were fixed. Also Khoury et al. [24] explored the suitability of ChatGPT for this purpose. They conducted a study where they first instructed ChatGPT to generate programs targeting certain vulnerabilities. Once ChatGPT provided the code, they analyzed it to identify any evident vulnerabilities. If vulnerabilities were present, they prompted ChatGPT to identify them and produce a safer version of the code. The study showed that ChatGPT could correct several vulnerabilities with further prompts, but the process still needs a lot of human guidance.
@@
Gong et al. [40] investigated the performance of LLMs in code generation and vulnerability detection. Their results show that LLMs often produce code with vulnerabilities, with 75% of the tested code classified insecure. The study found that LLMs generate insecure code due to limited risk awareness and struggle with high false positive rates in detecting their own vulnerabilities. While advanced LLMs can repair up to 60% of insecure code from other LLMs, they perform poorly on their own code due to so-called self-repair blind spots. To address these issues, the researchers developed a tool that utilizes LLMs in an iterative repair process, supported by semantic analysis, significantly enhancing the security of LLM-generated code.
VII-C Fine-tuning for Vulnerability Detection and Fixing

### v3

@@ 5.4. Factors Influencing LLMs Performance
Despite the remarkable capabilities of LLMs in code analysis tasks, their effectiveness is influenced by various factors, including the prompts used (further explored in Section 8), and the state of the code being analyzed.
@@
One issue arises when resolved vulnerabilities remain in the code as comments or annotations left by developers. LLMs may interpret these resolved issues as existing vulnerabilities, reducing their effectiveness in detecting real security vulnerabilities. This highlights the need for standardized commenting practices to improve LLMs performance. Additionally, when handling longer code snippets, LLMs often struggle to identify smaller, more subtle security defects (Yu et al., 2024).
@@
Several other challenges affect LLMs performance in detecting vulnerabilities. These include difficulty in understanding developers’ intent, misinterpretation of code comments, and inability to execute code. Chen et al. (Chen et al., 2024a) found that these challenges contributed to the high rate of false positives obtained when experimenting with ChatGPT as a vulnerability detection tool for smart contracts.
@@
Some authors have focused their research on identifying and addressing the factors that influence LLMs performance in code analysis tasks, aiming to enhance their accuracy and reliability. Specifically, Wang et al. (Wang et al., 2024) investigated the impact of naming on LLMs code analysis. In their experiment, they created datasets containing code with nonsensical or misleading variables, methods, and function names. Using CodeBERT for the analysis, they demonstrated that naming significantly affects LLMs ability to detect vulnerabilities, indicating that LLMs rely on clear and meaningful names.
@@
Fang et al. (Fang et al., 2024) identified a gap in research, concerning how effective LLMs are at analyzing confusing, obfuscated, or poorly readable code. The authors found that GPT models were more likely than smaller models to generate accurate and detailed explanations for input code. On the other hand, obfuscated code impacted the ability of LLMs to generate explanations. GPT models had a drop in security analysis accuracy, while smaller models were unable to handle obfuscated code. In conclusion, GPT models excel in code analysis but struggle with obfuscated code, indicating that clarity of the code significantly impacts LLMs analysis performance.
5.5. LLMs Fine-tuning for Vulnerability Detection
@@
Figure 3. Fine-tuning LLM with the security-specific dataset.
@@
In their work, Shestov et al. (Shestov et al., 2024) fine-tuned the LLM model WizardCoder (Luo et al., 2024) to improve it in vulnerability detection tasks for Java code. Their first contribution involved modifying the training process, which resulted in faster training times without compromising the model performance. The fine-tuned WizardCoder outperformed models such as CodeBERT, demonstrating the potential of fine-tuning LLMs for specialized source code analysis tasks. The authors also noted that vulnerability detection faces a fundamental challenge of class imbalance, with far fewer positive cases (i.e., the presence of vulnerabilities) than negative cases. By applying a combination of focal loss and sample weighting, the authors showed that models can achieve notable performance, even when dealing with such imbalances.
@@
After evaluating LLMs for vulnerability detection, the authors of (Liu et al., 2024a) fine-tuned one of the models that performed poorly on vulnerability detection tasks, Meta-Llama-3-8B-Intruct, to verify the impact of fine-tuning on the domain-specific performance. The results showed that fine-tuning can have a positive impact on vulnerability detection tasks, but challenges in identifying exact vulnerabilities remain. Moreover, fine-tuning was also explored by Ding et al. (Ding et al., 2025). Despite applying advanced techniques, such as class weights and contrastive learning, LLMs struggled with detecting vulnerabilities.
@@
Zhang et al. (Zhang et al., 2024d) introduced two fine-tuning approaches for vulnerability localization: discriminative fine-tuning and generative fine-tuning. The discriminative fine-tuning approach treats vulnerability localization as a sequence labeling task, classifying each code statement as either vulnerable or not. In contrast, generative fine-tuning trains models to generate both the line numbers and the corresponding vulnerable statements within a function. Currently, although generative fine-tuning appears promising, it is outperformed by discriminative fine-tuning, which significantly enhances model performance.
@@
Alqarni et al. (Alqarni and Azim, 2022) conducted experimental research on BERT with the aim to improve its ability on vulnerability detection tasks. Although BERT was not originally designed for this domain, they improved its performance by integrating additional layers into a new model, and retraining it using a GitHub vulnerability dataset. The BERT base models (two versions) use token, segment, and position embeddings to understand context, masking 15% of tokens to predict them during training. This new model, which integrates with BERT and shares both embeddings and attention mechanisms, achieved a notable test accuracy of 99.30%.
@@
Basharat et al. (Basharat and Omar, 2024) developed a method for detecting vulnerabilities in code using LLMs, by combining pattern-exploiting training and cloze-style queries to fine-tune models. Following the model of cloze tests, they hid parts of the code and asked the model to predict the missing parts. By performing this task on datasets containing vulnerable code, the model learns to recognize patterns that indicate potential security weaknesses. Results indicate that the method, especially using the RoBERTa model, can effectively identify vulnerabilities in different programming languages, such as C/C++.
@@
Guo et al. (Guo et al., 2024) compared 6 LLMs trained for vulnerability detection against 6 general-purpose LLMs, 3 of which were further fine-tuned on a dataset they compiled. Their findings demonstrate that fine-tuning improves the performance of smaller models, enabling them to outperform larger, general-purpose models in specific contexts. Fine-tuned models achieved remarkable results, with CodeLlama-7b-fine-tuned obtaining an F1-score of 97%, compared to the best performance of around 51% obtained by non-fine-tuned models. The study revealed that, while fine-tuning smaller models can improve their performance, they are limited to the exact tasks for which they were trained. Additionally, the research underscores issues with current datasets, including labeling errors, which can adversely affect model training and performance.
6. LLMs for Vulnerability Fixes
@@
Developers often introduce security vulnerabilities in their code, which, if exploited, can compromise system and application security. Addressing these vulnerabilities presents significant challenges, as manual mitigation is time-consuming, prone to errors, and requires experienced developers with specialized knowledge.
While prior work has shown that LLMs can be effective at detecting vulnerabilities, their capabilities do not end there. Beyond identifying vulnerabilities, LLMs can assist in identifying the root causes of these issues and providing the right fixes, significantly reducing developers’ workloads (Pearce et al., 2023; Zhang et al., 2024e). This section explores the usage of LLMs for fixing vulnerabilities, as well as the possibility of improving this usage through fine-tuning processes.
6.1. Uses of LLMs for Fixing Vulnerabilities
@@
Several studies focused on LLMs abilities to fix security vulnerabilities in code. For instance, Nong et al. (Nong et al., 2024b) used the power of LLMs for vulnerability fixing without needing specific test inputs or training/fine-tuning. To achieve this, they developed LLMPatch, an automated LLM-based fixing system. They evaluated LLMPatch on real-world vulnerable code, with a focus on zero-day vulnerabilities. LLMPatch was able to fix vulnerabilities and successfully surpassed traditional fixing strategies based on DL (VulRepair (Fu et al., 2022)) and ASTs (Getafix (Bader et al., 2019)).
@@
Berabi et al. (Berabi et al., 2024) proposed a method for enhancing LLMs abilities to fix code vulnerabilities. To increase the model performance, their approach combines static code analysis with code reduction, a technique that involves selecting only the code that contains a defect and the relevant context for interpreting it, and feeding both to the LLM. Results showed that this approach improved the efficiency of the model, by allowing it to understand and focus only on the sections that require fixing.
@@
Zhang et al. (Zhang et al., 2024e) demonstrated the ability of ChatGPT-4 and Claude to fix memory corruption vulnerabilities in real-world C/C++ code. LLMs can effectively address localized issues, such as memory leaks (CWE-401), but struggle with complex issues involving external dependencies, precise context, and intricate code structures. Furthermore, the authors highlighted that providing more context, such as external functions, data structures, and explanations of code logic, showed a positive effect on LLMs performance.
@@
Pearce et al. (Pearce et al., 2023) explored how well LLMs can fix software vulnerabilities in a zero-shot context. They evaluated the performance of 5 commercially available, black-box, “off-the-shelf” LLMs, as well as an open-source model, and their own locally-trained model on their dataset. They discovered that, while these models could successfully repair synthetic and hand-crafted vulnerabilities, they faced challenges with real-world scenarios, due to the complexity of prompt design and natural language variations.
6.2. LLMs Fine-tuning for Fixing Vulnerabilities
@@
Several researchers explored fine-tuning as a way to enhance LLMs effectiveness in fixing vulnerabilities. For instance, De et al. (de Fitero-Dominguez et al., 2024) fine-tuned two LLMs, Code Llama and Mistral, and compared them with traditional DL-based approaches, including VulRepair, VRepair (Chen et al., 2022), and VulMaster (Zhou et al., 2024b). The study also highlighted the problem of data overlap in previous research, revealing that models trained on overlapping datasets showed inflated performance metrics. To address this, their evaluation involved two datasets: an original dataset, used in other papers, and a refined dataset, which was curated by the authors to avoid overlaps between training and test data. This approach helped ensure a more accurate assessment of the model performance on new, unseen data. The results showed that the proposed method offered significant improvements in terms of code modifications efficiency and accuracy, compared to the DL-based methods.
@@
Wu et al. (Wu et al., 2023) evaluated the performance of different LLMs in fixing Java vulnerabilities, with and without fine-tuning, on two proposed benchmarks. Across 5 non-fine-tuned LLMs, the best-performing model, Codex, fixed 20.4% of vulnerabilities. Concerning the fine-tuning experiment, the authors could fine-tune 4 models out of 5, since OpenAI does not provide a public API for fine-tuning Codex. Among the fine-tuned models, the best performer (i.e., fine-tuned InCoder), fixed only 18% of vulnerabilities, a small improvement of 8% over the non-fine-tuned version. Besides, these 4 fine-tuned models succeeded only with simple fixes, such as deleting statements, renaming variables, or renaming methods. The models struggled with more complex vulnerabilities, such as CWE-325: Missing Cryptographic Step, and CWE-444: HTTP Request/Response Smuggling.
7. LLMs for Detecting and Fixing Vulnerabilities in a Unified Process
@@
Ensuring the security of increasingly complex software systems requires novel approaches that can keep pace with evolving threats. LLMs represent a promising tool for detecting and fixing software vulnerabilities (Le et al., 2024).
@@
In Section 5 and Section 6, we discussed using LLMs as tools for handling vulnerability detection and vulnerability fixing separately. In this section, we explore an integrated approach where LLMs are utilized to detect and fix security vulnerabilities in a unified process. Additionally, it examines the capability of LLMs to identify vulnerabilities in the code they generate, and automatically resolve them. Last, the section covers fine-tuning for enhancing LLMs performance, in a unified pipeline for vulnerability detection and fixing.
7.1. LLMs for Detecting and Fixing Vulnerabilities
@@
Several studies have focused on both detecting and fixing software vulnerabilities using LLMs. For instance, Noever et al. (Noever, 2023) evaluated the capability of GPT-4 to detect software vulnerabilities and suggest fixes for such vulnerabilities. GPT-4 demonstrated a strong ability to detect and fix vulnerabilities, achieving a 90% reduction in vulnerabilities with only an 11% increase in lines of code, while maintaining low rates of false positives.
@@
Le et al. (Le et al., 2024) investigated the efficiency of ChatGPT and Bard in the detection and fixing of security vulnerabilities, focusing on JavaScript code. The results showed that ChatGPT outperformed Bard with a marginal higher accuracy of 71.66%, versus 68.33%.
7.2. LLMs for Detecting and Fixing Vulnerabilities in Self-produced Code
@@
As we discussed in Section 4, code generated by LLMs may contain various vulnerabilities. However, researchers demonstrated that it is possible to re-prompt LLMs to identify and fix those vulnerabilities. This approach has demonstrated that iterative prompting can lead to significant improvements in code security.
@@
The research conducted by Liu et al. (Liu et al., 2024c) has shown promising results in addressing security vulnerabilities using ChatGPT, with over 89% of vulnerabilities successfully mitigated during a multi-round fixing process. In this approach, Liu et al. interacted with ChatGPT over several rounds, giving feedback on the provided code, and asking for improvements until the vulnerabilities were fixed. Khoury et al. (Khoury et al., 2023) conducted a study where they first instructed ChatGPT to generate programs targeting certain vulnerabilities. Once ChatGPT provided the code, they analyzed it to identify any evident vulnerabilities. If vulnerabilities were present, they prompted ChatGPT to identify them and produce a safer version of the code. The study showed that ChatGPT could correct several vulnerabilities with further prompts, but the process still needs a lot of human guidance.
@@
Gong et al. (Gong et al., 2024) investigated the performance of LLMs in code generation and vulnerability detection. Their results show that LLMs often produce code with vulnerabilities, with 75% of the tested code classified insecure. The study found that LLMs generate insecure code due to limited risk awareness and struggle with high false positive rates in detecting their own vulnerabilities. While advanced LLMs can fix up to 60% of insecure code from other LLMs, they perform poorly on their own code due to so-called self-repair blind spots. To address these issues, the researchers developed a tool that utilizes LLMs in an iterative fix process, supported by semantic analysis, significantly enhancing the security of LLM-generated code.
7.3. Fine-tuning for Vulnerability Detection and Fixing